OpenStudy (anonymous):

yes or no if the question is correct 1) 5(3v + 5)=70>>.v=3 yes or no 2) -5(1-3m)=55>>>m=1 yes or no 3) -5(-3n - 2)=55>>>N=3 yes or no 4) 5(4a - 4)= -120>>a=4 yes or no

jimthompson5910 (jim_thompson5910):

1) is correct

jimthompson5910 (jim_thompson5910):

2) is incorrect

jimthompson5910 (jim_thompson5910):

3) is correct

jimthompson5910 (jim_thompson5910):

4) is incorrect

jimthompson5910 (jim_thompson5910):

2) -5(1-3m)=55 -5+15m=55 15m = 55+5 15m = 60 m = 60/15 m = 4

OpenStudy (anonymous):

whats 3 and 4 then

jimthompson5910 (jim_thompson5910):

3 is correct

jimthompson5910 (jim_thompson5910):

4) 5(4a - 4)= -120 20a - 20 = -120 20a = -120 + 20 20a = -100 a = -100/20 a = -5